# Roughening dynamics of interfaces in two-dimensional quantum matter

### v1

@@ Appendix C Kink operator
Figure S6: a) Time evolution of the bare, i.e. unmodified kink operator in the 2D TFIM. b) Time evolution of the modified kink operator for transverse fields $g/J=1,1.25,1.5$, for both the 2D TFIM and 1D effective model. The plot shows unphysical growth of the modified kink operator for $g/J>1$, signaling the breakdown of the assumption that bulk and interface effects can be separated. Plots c) and d) show the modified, time averaged kink operator as a function of c) the angle $\alpha$ and d) the distance $l$. The distance does not have much influence on how well the effective and full model agree, which is why throughout the work we chose the largest distance possible for our lattice, i.e. $l=8$. This is different for the angle where we observe larger deviations for the largest transverse field $g/J=1$. The smallest angle which showed the smallest deviations across all values of the transverse fields considered is $\alpha=1$, which we used throughout this work.
@@
Appendix D Exact Solution of the Classical Limit and Approximations
@@
Since the distance between two solutions vanishes as $1/M$ in the large $M$ limit, we expect $|\theta_{n_{\star}}-\theta_{\star}|\approx\delta/M$, where $\delta$ is a constant of order and independent of $\alpha$. Expanding $c_{n_{\star}}$ for small $\delta/M$ and keeping only the dominant part, we arrive at
@@
showing the expected divergent behavior with $M$.
@@
Appendix E Quantum Monte Carlo
The data points in the phase diagram of Fig. 2d) are obtained using the QMC loop-algorithm [36, 37] provided in the ALPS package [56].
@@ Appendix E Quantum Monte Carlo
The points of the red line correspond to the temperature associated with the initial domain wall state on an $8\times 8$ lattice at a transverse field $g$, i.e. the temperature $T_{\mathrm{cross}}$ s.t. the energy $E=\mathrm{Tr}\ H\rho_{\mathrm{Gibbs}}$ of the Gibbs state $\rho_{\mathrm{Gibbs}}=\exp{(-H/T_{\mathrm{cross}})}$ corresponds to the energy of the initial state $E_{\mathrm{init}}$, see Fig. S8 b.
All data points are obtained using $5\cdot 10^{4}$ samples, after $1\cdot 10^{4}$ thermalization steps are performed, these values where chosen such that the estimated error on all the observables considered here lies below $1\%$.

### v2

@@ Appendix D Kink operator
Figure S6: a) Time evolution of the bare, i.e. unmodified kink operator in the 2D TFIM. b) Time evolution of the modified kink operator for transverse fields $g/J=1,1.25,1.5$, for both the 2D TFIM and 1D effective model. The plot shows unphysical growth of the modified kink operator for $g/J>1$, signaling the breakdown of the assumption that bulk and interface effects can be separated. Plots c) and d) show the modified, kink operator, averaged over the interval $tJ\in[20,100]$, as a function of c) the angle $\alpha$ and d) the distance $l$. The distance does not have much influence on how well the effective and full model agree, which is why throughout the work we chose the largest distance possible for our lattice, i.e. $l=8$. This is different for the angle where we observe larger deviations for the largest transverse field $g/J=1$. The smallest angle which showed the smallest deviations across all values of the transverse fields considered is $\alpha=1$, which we used throughout this work.
@@
Results for the kink operator on a 16 by 16 lattice are provided in Figure S7, in a similar fashion to what has been presented in the main text for an 8 by 8 lattice. Due to the larger overall entanglement, the simulations only provide reliable results up to a timescale of $t_{\mathrm{max}}=30$.
Figure S7 a) shows the domain wall length contribution coming from horizontally aligned interface lengths, normalized by the length of the lattice in x-direction. Once again this confirms that up to $g/J=1$, the number of domain walls per column does not scale extensively with the vertical lattice size. The agreement between the time evolution of the modified kink operator calculated in the full 2D and effective models is shown in Figure S7 b).
We want to emphasize that for that comparison we did not utilize the running mean of the modified kink operator as has been done in the main text for the 8 by 8 system. This is because compared to the 8 by 8 system, the kink operator calculated in the 16 by 16 system does not show strong temporal fluctuations, allowing for a direct comparison.
The agreement between the two models is made even more apparent when comparing the late time plateau values of the respective kink operators, obtained by taking the mean points within the region $tJ\in[5,30]$, see Figure S7 c).
Figure S7: Kink operator for a 16 by 16 system, evolved up to a time of $t_{\max}=30$. a) Time evolution of the horizontal contribution to the domain wall length divided by the horizontal lattice dimension. The plot shows that the number of horizontal domain walls in each column does not scale extensively with the vertical system size, confirming the validity of the effective model in that regime. (b) Comparison between the effective (dashed lines/crosses) and the full model (solid lines/circles) for the time evolution of the kink operator. For the full model, we calculate the modified kink operator $K_{\mathrm{M}}=\langle K\rangle/\langle K_{\mathrm{bulk}}\rangle$. We don’t show data for transverse fields beyond $g/J=1$, as bulk and interface contributions become increasingly difficult to disentangle. (c) Late-time averages over the interval $tJ\in[5,30]$ of the kink operators shown in (b) as functions of the transverse field. The values agree well up to $g/J=1$, once again confirming the quantitative predictive power of the effective in that regime.
Appendix E Exact Solution of the Classical Limit and Approximations
@@
Since the distance between two solutions vanishes as $1/M$ in the large $M$ limit, we expect $|\theta_{n_{\star}}-\theta_{\star}|\approx\delta/M$, where $\delta$ is a constant of order one and independent of $\alpha$. Expanding $c_{n_{\star}}$ for small $\delta/M$ and keeping only the dominant part, we arrive at
@@
which diverges as expected for $M\to\infty$.
@@
Appendix F Thermalization and level spacing statistics
In the main text, we implicitly assume eventual equilibration of the interface initial condition to a homogeneous spin state. To support this claim, which is not trivial and still under debate, we provide an analysis of the level spacing statistics of the exact eigenenergy spectrum $E_{i}$ for a 4 by 5 lattice with PBC in x-direction and OBC in y-direction, analogous to what has been done in the main text for larger system sizes. The level spacing statistics provides a way to determine wether a system is chaotic or integrable, based on results from random matrix theory. An integrable system is characterized by its level spacing statistics following a Poisson distribution, while for a chaotic system it follows the one of a Gaussian orthogonal ensemble (GOE). In order to evaluate the level spacing statistics reliably, one has to first unfold the spectrum, i.e. separate the level spacing from overall changes in the energy scale. This is done via the relation
$s_{n}=(E_{n+1}-E_{n})\frac{\partial n(E)}{\partial E}|_{E=E_{n}},$ (S14)
where $n(E)$ is the staircase or cumulative spectral function, returning the number of levels below or at the energy $E$ [43].
In Figure S9 a) we show results of the unfolded spectrum for two different transverse fields $g/J=0.15,0.75$. The level spacing of the larger field clearly resembles the one of a GOE, indicating a chaotic system, which is true also for the smaller transverse field, even though the peak of the distribution starts to travel towards zero level spacing. To summarize, neither of the cases follows the level spacing distribution expected from an integrable system.
The mean over the distribution of ratios
$\tilde{r}_{n}=\frac{\min(d_{n},d_{n-1})}{\max(d_{n},d_{n-1})},$ (S15)
with $d_{n}=E_{n}-E_{n-1}$, assumes the value $\langle\tilde{r}_{n}\rangle_{\mathrm{Poisson}}\approx 0.38629$ or $\langle\tilde{r}_{n}\rangle_{\mathrm{GOE}}\approx 0.60266$ if the underlying distribution is of Poissonian or GOE type, respectively [44]. Figure S9 b) presents this quantity as a function of the transverse field, showing that in the considered system of finite size the underlying level spacing statistics closely follows the chaotic one, with minimal deviations towards smaller transverse fields. Hence, we expect eventual thermalization of the system in the non-equilibrium dynamics considered in the main text for all system sizes considered.
Figure S9: a) Two examples of the unfolded level spacings at different $g/J$. The one corresponding to $g/J=0.75$ clearly agrees with the one predicted from random matrix theory for chaotic systems, while for the smaller $g/J=0.15$ it starts exhibiting features from both, i.e. the mean level spacing moves towards zero – however the characteristic peak at $s_{n}=0$ is still missing, distinguishing it from a pure Poisson statistics. b) Mean level-spacing distribution ratio of a 4 by 5 system for different transverse fields. Across all transverse fields considered, the ratio corresponds to a GOE distribution and hence an chaotic system with slight deviations from that expected value at the smallest fields.
Appendix G Quantum Monte Carlo
The data points in the phase diagram of Fig. 2d) are obtained using the QMC loop-algorithm [40, 41] provided in the ALPS package [60] for the critical temperatures (blue line) and using the QMC package written in Rust [64] for the effective temperature of the initial state (red line).
@@ Appendix G Quantum Monte Carlo
The points of the red line correspond to the temperature associated with the initial domain wall state on an $8\times 8$ lattice at a transverse field $g$, i.e. the temperature $T_{\mathrm{cross}}$ s.t. the energy $E=\mathrm{Tr}\ H\rho_{\mathrm{Gibbs}}$ of the Gibbs state $\rho_{\mathrm{Gibbs}}=\exp{(-H/T_{\mathrm{cross}})}$ corresponds to the energy of the initial state $E_{\mathrm{init}}$, see Fig. S10 b. Note, that in order to simulate the conditions for a domain wall state and get the correct effective temperature, we employ periodic boundary conditions on one edge and anti-periodic boundary conditions on the other within the QMC simulation.
Data points are obtained using $5\cdot 10^{4}$/$1\cdot 10^{5}$ samples for the critical/actual temperatures, after $1\cdot 10^{4}$ thermalization steps are performed. These values where chosen such that the estimated error on all the observables considered here lies below $1\%$.